OpenStudy (butterflydreamer):

CALCULATIONS - gravimetric analysis (please check my answer ^_^ The data from an experiment to determine the % (by mass) of sulfate in lawn fertiliser are provided. Mass of fertiliser sample =10.000 g Mass of filter paper= 0.0250 g Mass of filter paper and dry BaSO4 = 0.485 g Calculate the percentage of sulfate in the fertiliser

OpenStudy (butterflydreamer):

Basically what i did was let y = mass of fertiliser sample Therefore y= 10 g Then we find the (formula mass of SO4) / (formula mass of BaSO4) = 96.07 / 233.37 = 0.4117 Mass (SO4) = 0.4117 * mass (precipitate) = 0.4117 * (0.485 - 0.0250) = x g % w/w = x/y * 100 = 1.89 % Is that correct? I feel totally off xD

OpenStudy (aaronq):

in "= 0.4117 * (0.485 - 0.0250) " why did you write 0.485 and not 0.495?

OpenStudy (butterflydreamer):

oh snap. I made a typo in the question. It should be 0.485g NOT 0.495

OpenStudy (aaronq):

it's good otherwise

OpenStudy (butterflydreamer):

thaaank you!